40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 5 15:28:38 2014, Steve, Update, SUS, HeNe laser test preparation beforeRSMqpdMoved.jpgafterSRMqpdMoved.jpg1103Ptest.pngND1onQPD.png
    Reply  Fri Sep 5 16:25:36 2014, Steve, Update, SUS, center ITMY oplev 
    Reply  Tue Sep 9 08:22:28 2014, Steve, Update, SUS, HeNe laser test preparation HeNecompared.png
       Reply  Tue Sep 9 16:59:32 2014, Steve, Update, SUS, HeNe laser test preparation 1103PqpdTranss.png
          Reply  Wed Sep 10 14:58:58 2014, Jenne, Update, SUS, OpLev test: New channels 
             Reply  Thu Sep 11 17:12:42 2014, Steve, Update, SUS, OpLev test: old SP qpd connected 
                Reply  Fri Sep 12 16:30:48 2014, Steve, Update, SUS, HeNe laser test preparation ND2.pngqpd222ND21103P.jpg
                   Reply  Wed Sep 17 13:41:00 2014, Jenne, Update, SUS, New calibrated channels for test QPD NewQPDcalibratedChans.png
                      Reply  Thu Sep 18 09:17:47 2014, Steve, Update, SUS, HeNe laser test  calibrated_16hrs.pngQPD222cal.jpg1103P40s40min.png
                         Reply  Mon Sep 22 10:14:07 2014, Steve, Update, SUS, HeNe laser test  1103Pmon3d.png
                            Reply  Tue Sep 23 08:34:14 2014, Steve, Update, SUS, HeNe laser test  tempDrop.png
                               Reply  Thu Sep 25 10:53:32 2014, Steve, Update, SUS, HeNe laser test  centeredDay6.png
          Reply  Wed Sep 10 15:31:16 2014, Steve, Update, SUS, HeNe laser test preparation SRMqpdisBack0nSRMsus.pngslowVSfast.png
    Reply  Wed Oct 15 17:09:49 2014, manasa, Update, General, Diode laser test preparation 
       Reply  Wed Oct 29 18:07:28 2014, manasa, Update, General, Diode laser test preparation 
Message ID: 10459     Entry time: Fri Sep 5 15:28:38 2014     Reply to this: 10461   10475   10610
Author: Steve 
Type: Update 
Category: SUS 
Subject: HeNe laser test preparation 

The SRM qpd was moved to accommodate the HeNe laser qualification test for LIGO Oplev use.

The qpd was saturating at 65,000 counts of 3 mW

ND1 filter lowering the power by 10 got rid of saturation. I epoxied an adapter ring to the qpd.

Atm3 was taken before saturation was realized with Koji's help.

Atm4 ND1 on SRM qpd. Now it is working and everything is moving.

 

Attachment 1: beforeRSMqpdMoved.jpg  246 kB  | Hide | Hide all | Show all
beforeRSMqpdMoved.jpg
Attachment 2: afterSRMqpdMoved.jpg  228 kB  | Show | Hide all | Show all
Attachment 3: 1103Ptest.png  177 kB  Uploaded Fri Sep 5 16:30:27 2014  | Hide | Hide all | Show all
1103Ptest.png
Attachment 4: ND1onQPD.png  381 kB  Uploaded Fri Sep 5 17:14:14 2014  | Show | Hide all | Show all
ELOG V3.1.3-